SUM FREQUENCY VIBRATIONAL SPECTROSCOPY STUDY OF THE IONIC LIQUID-TITANIUM DIOX-IDE INTERFACE

\underline{CESAR ALIAGA} and STEVEN BALDELLI, Department of Chemistry, University of Houston, Houston, TX 77004.

Sum frequency Generation Spectroscopy is performed at the ionic liquid-titanium dioxide interface. Two pure room temperature ionic liquids are studied: 1-butyl-3-methylimidazolium methanesulfonate ([BMIM][MS]), and 1-butyl-3-methylimidazolium Dicyanamide ([BMIM][DCA]). Vibrational modes pertaining to the cation and the anion are detected, molecular orientation calculations are performed, and a structure of the ordering of the molecular species at the interface is proposed.